### v4

@@ Abstract
The proliferation of 100B+ parameter Large Language Models (LLMs) with 100k+ context length support have resulted in increasing demands for on-chip memory to support large KV caches. Techniques such as StreamingLLM xiao2024efficientstreaminglanguagemodels and SnapKV li2024snapkvllmknowslooking demonstrate how to control KV cache size while maintaining model accuracy. Yet, these techniques are not commonly used within industrial deployments using frameworks like vLLM or SGLang. The reason is twofold: on one hand, the static graphs and continuous batching methodology employed by these frameworks kwon2023efficientmemorymanagementlarge; zheng2024sglangefficientexecutionstructured make it difficult to admit modifications to the standard multi-head attention algorithm vllmstreamingllmsupport2023, while on the other hand, the accuracy implications of such techniques on modern instruction-following and reasoning models are not well understood, obfuscating the need for implementing these techniques. In this paper, we explore these accuracy implications on Llama-3.1-8B-Instruct and DeepSeek-R1, and develop SnapStream, a KV cache compression method that can be deployed at scale. We demonstrate the efficacy of SnapStream in a 16-way tensor-parallel deployment of DeepSeek-671B on SambaNova SN40L accelerators running at 128k context length and up to 1832 tokens per second in a real production setting. SnapStream enables $4\times$ improved on-chip memory usage and introduces minimal accuracy degradation on LongBench-v2, AIME24 and LiveCodeBench. To the best of our knowledge, this is the first implementation of sparse KV attention techniques deployed in a production inference system with static graphs and continuous batching.
@@ 1 Introduction
In response, practitioners have developed a variety of architectural modifications and compression methods to alleviate this memory pressure. Multi-query attention ainslie2023gqatraininggeneralizedmultiquery, multi-head latent attention deepseekai2024deepseekv2strongeconomicalefficient and sliding window attention gemmateam2024gemma2improvingopen each modify the attention operation to reduce memory requirements of the attention head, hidden and sequence dimension, respectively. However, these methods generally involve training a model from scratch with the proposed attention alternatives, reducing the applicability for existing models. KV cache compression methods, such as H2zhang2023h2oheavyhitteroracleefficient and SnapKV li2024snapkvllmknowslooking, take advantage of the high sparsity of the attention matrix zhang2025spargeattentionaccuratetrainingfreesparse to evict tokens with low attention scores from the KV cache. Such methods are largely training-free and can reduce KV cache memory by up to 92% with negligible decrease in benchmark scores.
However, when applying these techniques to modern LLM production deployments, the following practical considerations emerge:
@@
  • We introduce SnapStream, an efficient KV cache compression method designed for LLM serving at scale with continuous batching deployments. We describe the structure of the compressed KV cache, and the modifications that must be made during prefill and decoding phases respectively to implement SnapStream.
  • We describe fused kernels and tensor sharding strategies for a static graph deployment of DeepSeek-R1 on 16 SN40L accelerators. We show how different mapping strategies for prefill and decoding optimize for different service level objectives (SLOs), and how SnapStream can be accommodated by adding fused kernels to the prefill execution schedule.
@@ 2.1 Continuous Batching
LLM serving is defined by two stages of computation: the prefill phase and the decode phase, each with their own service level objectives (SLOs): time to first token (TTFT) and time per output token (TPOT), respectively. In prefill, the LLM processes a user prompt of sequence length $L$ to generate a new token $x_{L+1}$ and KV cache tensors of length $L$. In this phase, the $O(L^{2})$ attention mechanism dominates latency, making the prefill phase compute bound. In the decoding phase, the model processes a KV cache of length $L$ and a single input token $x_{L+1}$ to generate the next token $x_{L+2}$ and the KV cache of length $L+1$. By re-using the key and value tensors in the KV cache, multi-head attention becomes $O(L)$ during decoding. With significantly lower computational intensity compared to the prefill phase, the decode phase is typically memory bound. In practice, compute-intensive long sequence prefill requests are often served with low batch size to minimize TTFT, while memory-intensive decoding requests are served with higher batch size to maximize hardware utilization.
Continuous batching (CB) is a popular LLM serving technique Yu2022OrcaAD; agrawal2023sarathiefficientllminference; kwon2023efficientmemorymanagementlarge in which batches of prefill and decoding requests may be interleaved on the same system, and the output of completed prefill requests can be added to decode batches to improve utilization. However, there is an inherent contention to the different SLOs for prefill and decoding, as indicated by the difference between their batching preferences. Disaggregated Continuous Batching (DCB) zhong2024distservedisaggregatingprefilldecoding proposes to resolve this contention by assigning separate resources to the prefill and decoding phases and only transferring KV caches between nodes. zhong2024distservedisaggregatingprefilldecoding show that DCB can sustain up to a $7.4\times$ request rate compared to the leading CB baselines on a 32-GPU, 4-node setup.
@@ 2.2 Reconfigurable Dataflow Units (RDUs)
Built on a dataflow architecture, the SN40L AI accelerator optimizes AI training and inference tasks. Prior works Prabhakar_2024; sn40l-hotchips describe the architecture in more detail. This section provides a brief overview of key architectural components. Each socket features four RDU cores, I/O blocks for connecting to DDR, HBM, Peer-to-Peer (P2P) links, and a host system (see Figure 2). A socket can perform 638 TFLOPs of BF16 operations and has 520 MB of on-chip SRAM. Each socket also interfaces with 64 GB of HBM memory providing 1.6 TB/s of peak bandwidth, and 1.5 TB of DDR memory providing over 100 GB/s of peak bandwidth. A Top-Level Network (TLN) connects tile components to the IO blocks. Each socket consists of 1040 Pattern Compute Units (PCUs), which provide systolic and streaming compute capabilities, and 1040 Pattern Memory Units (PMUs) which provide program-managed on-chip memory for storing tensors. Collectively, a PCU and PMU block is known as a Fused Compute Memory Unit (FCMU). Additionally, the Address Generation and Coalescing Units (AGCUs) serve as the interface to external memory and other sockets.
@@ 2.3 LLM Serving with Static Graphs
The SN40L compiler operates primarily on static graphs, where all operations are executed on fixed size tensors. Specifically, for an LLM, all operations are executed with a fixed sequence length $L_{\text{max}}$, so that given an input token sequence $x_{1},...,x_{L}$ with $L\leq L_{\text{max}}$, we operate on the padded sequence $x_{1},...,x_{L},x_{\text{pad}:L+1},...,x_{\text{pad}:L_{\text{max}}}$. For the $i$th hidden layer and $j$th attention head, executing the prefill graph generates key and value sequences $k^{i,j}_{1},...,k^{i,j}_{L},k^{i,j}_{\text{pad}:L+1},...,k^{i,j}_{\text{pad}:L_{\text{max}}}$ and $v^{i,j}_{1},...,v^{i,j}_{L},v^{i,j}_{\text{pad}:L+1},...,v^{i,j}_{\text{pad}:L_{\text{max}}}$.  [^1] For brevity, in future references to these variables we will omit the layer and head indices $i,j$. Similarly, we will only describe operations on the key values $k$, as our treatment of the value tensors $v$ is mostly identical. The prefill graph also generates a single new token $x_{L+1}$. Executing the decode graph immediately after prefill generates new key and value tensors $k_{L+1}$ and $v_{L+1}$ that are used to update the KV cache in-place (e.g. $k_{1},...,k_{L},k_{L+1},k_{\text{pad}:L+2}...,k_{\text{pad}:L_{\text{max}}}$) and generates one more new token $x_{L+2}$. The new token and the updated KV cache are used as input for another call to the decode graph. This process is repeated until the EOS token or $x_{L_{\text{max}}}$ is generated. See Figure 3 (a) and (c) for a visualization of this process.
@@ 2.4 SnapKV
SnapKV is a KV cache compression method that mitigates the lossy compression seen in prior compression schemes by using a pooling-based clustering mechanism. The pooling layer helps to retain critical token clusters with higher attention weights, thus retaining the completeness of information. This clustering is particularly beneficial in retrieval tasks like needle-in-a-haystack, as we later show in the Retrieval subset of RULER in Table 1. Formally, given a prompt with length $L=L_{\text{prefix}}+L_{\text{obs}}$ divided into a prefix and an observation window, we obtain the SnapKV index set $I$ by computing:
$C$ $=\sum_{i=0}^{L_{\text{obs}}}W[:,:,i,:]$ (1)
$I$ $=\text{Top}_{K}(\text{avgpool}(C))$ (2)
where $W\in\mathbb{R}^{B\times N\times L_{\text{obs}}\times L_{\text{prefix}}}$ is the softmax-normalized attention distribution of tokens in the observation window attending to prefix tokens, with batch size $B$ and number of attention heads $N$.
@@ 2.5 StreamingLLM
StreamingLLM xiao2024efficientstreaminglanguagemodels is a technique that enables LLMs to process sequences beyond their fixed context windows, effectively supporting unbounded streaming inputs. Conventional LLMs are limited by a fixed attention window, which constrains their ability to maintain coherence over long sequences. StreamingLLM addresses this by retaining a small set of attention sink tokens (the initial prompt tokens that consistently receive high attention across decoding steps) together with the most recent tokens. These tokens act as stable anchors that preserve long-term context, allowing the model to maintain continuity without storing the entire KV cache.
The core mechanism of StreamingLLM is a compact KV cache strategy that combines these sink and recent tokens to form a reduced attention context. During decoding, the model attends primarily to this subset, substantially lowering memory usage while preserving model accuracy.
@@ 3 Methodology
SnapStream is built on top of a static graph CB deployment by applying SnapKV to prefill KV caches and decoding with StreamingLLM. Following xiao2024efficientstreaminglanguagemodels, given an input sequence of length $L$, during prefill we extract the first $L_{\text{sink}}$ tokens and the last $L_{\text{recent}}$ tokens from the full KV cache. From the remaining $L-(L_{\text{sink}}+L_{\text{recent}})$ tokens, we apply SnapKV compression to extract the top $K$ tokens. We concatenate all three together to obtain a prefill KV cache of length $L_{\text{snapstream}}=L_{\text{sink}}+L_{\text{recent}}+K$. During decoding, we decode with a fixed KV cache size of $L_{\text{snapstream}}$ treating the middle $L_{\text{recent}}$ tokens as the rotating KV cache described in xiao2024efficientstreaminglanguagemodels. See Figure 3(b) and (d) for a visualization of SnapStream prefill and decoding, respectively. In the following sections, we describe the layout of the SnapStream KV cache and the operation of SnapStream prefill and decoding in a production static-graph execution framework.
@@ 3.1.1 Case 1: $L<L_{sr}$
In this case, the key tensor cache is constructed during prefill as $k_{1},...,k_{L},k_{\text{pad}:L+1},..,k_{\text{pad}:L_{sr}},k_{\text{topk}:1},..,k_{\text{topk}:K}$. This structure is no different than that described in Section 2.3, as the Top-$K$ tokens $x_{\text{topk}:\cdot}$ are themselves padding tensors in this case. Accordingly, during decoding, the KV cache is updated in-place in the same manner as previously described.
@@ 3.1.2 Case 2: $L\geq L_{sr}$
In this case, the key cache is constructed as $k_{1},...,k_{L_{\text{sink}}},k_{L-L_{\text{recent}}},...,k_{L},k_{\text{topk}:1},...,k_{\text{topk}:K}$. During prefill, we remove the middle $L-L_{sr}$ tokens from the KV cache, and extract the top $K$ key and value tensors from these tokens to popuate $k_{\text{topk}:1},...,k_{\text{topk}:K}$. During decoding, we maintain the same KV cache size by evicting the least recently used key tensor, $k_{L-L_{\text{recent}}}$, and replacing it with the new key tensor $k_{L+1}$. Note that during decoding, the sink tensors and top-$K$ tensors are never modified.
@@ 3.2 Prefill
In order to reduce memory usage of the decoding stage, SnapStream applies SnapKV compression during prefill to compress an output KV cache along the sequence dimension. Adherent to the KV cache structure described in Section 3.1, we only apply SnapKV to compress the removed tokens $k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}$. The sink token KV cache, recent token KV cache, and compressed SnapKV cache are then concatenated and sent to the decoding node.
@@
  • We apply StreamingLLM after any positional encodings are applied to query, key, and value projections. This means that we do not apply positional encodings to cached KVs during decoding and use the original relative positional encodings. Although this may affect the ability of the model to generate beyond its trained attention window size, as claimed in xiao2024efficientstreaminglanguagemodels Section 3.2, we believe that this change is well-motivated, as most modern open models are now trained to handle context lengths up to 128k grattafiori2024llama3herdmodels; deepseekai2025deepseekv3technicalreport; yang2025qwen3technicalreport.

### v5

@@ Abstract
The proliferation of 100B+ parameter Large Language Models (LLMs) with 100k+ context length support have resulted in increasing demands for on-chip memory to support large KV caches. Techniques such as StreamingLLM xiao2024efficientstreaminglanguagemodels and SnapKV li2024snapkvllmknowslooking demonstrate how to control KV cache size while maintaining model accuracy. Yet, these techniques are not commonly used within industrial deployments using frameworks like vLLM or SGLang. The reason is twofold: on one hand, the static graphs and continuous batching methodology employed by these frameworks kwon2023efficientmemorymanagementlarge; zheng2024sglangefficientexecutionstructured make it difficult to admit modifications to the standard multi-head attention algorithm vllmstreamingllmsupport2023, while on the other hand, the accuracy implications of such techniques on modern instruction-following and reasoning models are not well understood, obfuscating the need for implementing them. In this paper, we explore these accuracy implications on Llama-3.1-8B-Instruct and DeepSeek-R1, and develop SnapStream, a KV cache compression method that can be deployed at scale. We demonstrate the efficacy of SnapStream in a 16-way tensor-parallel deployment of DeepSeek-671B on SambaNova SN40L accelerators running at 128k context length and up to 1832 tokens per second in a real production setting. SnapStream enables $4\times$ improved on-chip memory usage and introduces minimal accuracy degradation on LongBench-v2, AIME24 and LiveCodeBench. To the best of our knowledge, this is the first implementation of sparse KV attention techniques deployed in a production inference system with static graphs and continuous batching.
@@ 1 Introduction
In response, practitioners have developed a variety of architectural modifications and compression methods. Multi-query attention ainslie2023gqatraininggeneralizedmultiquery, multi-head latent attention deepseekai2024deepseekv2strongeconomicalefficient and sliding window attention gemmateam2024gemma2improvingopen each modify the attention operation to reduce memory requirements of the attention head, hidden and sequence dimensions respectively. However, these methods generally involve training a model from scratch with the proposed attention alternatives, reducing their applicability to existing models. KV cache compression methods, such as H2zhang2023h2oheavyhitteroracleefficient and SnapKV li2024snapkvllmknowslooking, take advantage of high sparsity levels in the attention matrix zhang2025spargeattentionaccuratetrainingfreesparse to evict tokens with low attention scores from the KV cache. Such methods are largely training-free and can reduce KV cache memory by up to 92% with negligible decrease in benchmark scores. However, when applying these techniques to modern LLM production deployments, the following practical considerations emerge:
@@
  • We introduce SnapStream, an efficient KV cache compression method designed for LLM serving at scale with continuous batching deployments. We describe the structure of the compressed KV cache, and the modifications that must be made to prefill and decoding phases respectively to implement SnapStream.
  • We describe kernel fusion and tensor sharding strategies for a static graph deployment of DeepSeek-R1 on a SN40L-16 node. We show how different mapping strategies for prefill and decoding optimize for different service level objectives (SLOs), and how SnapStream can be accommodated by adding fused kernels to the prefill execution schedule.
@@ 2.1 Continuous Batching
LLM serving is defined by two stages of computation: the prefill phase and the decode phase, each with their own service level objectives (SLOs). These SLOs are time to first token (TTFT) and time per output token (TPOT) for prefill and decoding respectively. In prefill, the LLM processes a user prompt of sequence length $L$ to generate a new token $x_{L+1}$ and KV cache tensors of length $L$. In this phase, the $O(L^{2})$ attention mechanism dominates latency, making the prefill phase compute bound. In the decoding phase, the model processes a KV cache of length $L$ and a single input token $x_{L+1}$ to generate the next token $x_{L+2}$ and the KV cache of length $L+1$. By re-using the key and value tensors in the KV cache, multi-head attention becomes $O(L)$ during decoding. With significantly lower computational intensity compared to the prefill phase, the decode phase is typically memory bound. In practice, compute-intensive long sequence prefill requests are often served with low batch size to minimize TTFT, while memory-intensive decoding requests are served with higher batch size to maximize hardware utilization.
Continuous batching (CB) is a popular LLM serving technique Yu2022OrcaAD; agrawal2023sarathiefficientllminference; kwon2023efficientmemorymanagementlarge in which batches of prefill and decoding requests may be interleaved on the same system, and the output of completed prefill requests can be added to decode batches to improve utilization. However, there is an inherent contention to the different SLOs for prefill and decoding, as indicated by the difference between their batching preferences. Disaggregated Continuous Batching (DCB) zhong2024distservedisaggregatingprefilldecoding proposes to resolve this contention by assigning separate hardware resources to the prefill and decoding phases and transferring KV caches from prefill to decoding nodes between phases. The authors show that DCB can sustain up to a $7.4\times$ request rate compared to the leading CB baselines on a 32 GPU, 4 node setup.
@@ 2.2 Reconfigurable Dataflow Units (RDUs)
Built on a dataflow architecture, the SN40L AI accelerator optimizes AI training and inference tasks. Prior works Prabhakar_2024; sn40l-hotchips describe the architecture in more detail. This section provides a brief overview of key architectural components. Each socket features four RDU cores, I/O blocks for connecting to DDR, high bandwidth memory (HBM), Peer-to-Peer (P2P) links, and a host system (see Figure 2). A socket can perform 638 TFLOPs of BF16 operations and has 520 MB of on-chip SRAM. Each socket also interfaces with 64 GB of HBM memory providing 1.6 TB/s of peak bandwidth, and 1.5 TB of DDR memory providing over 100 GB/s of peak bandwidth. A Top-Level Network (TLN) connects tile components to the IO blocks. Each socket consists of 1040 Pattern Compute Units (PCUs), which provide systolic and streaming compute capabilities, and 1040 Pattern Memory Units (PMUs) which provide program-managed on-chip memory for storing tensors. Collectively, a PCU and PMU block is known as a Fused Compute Memory Unit (FCMU). Additionally, the Address Generation and Coalescing Units (AGCUs) serve as the interface to external memory and other sockets.
@@ 2.3 LLM Serving with Static Graphs
The SN40L compiler operates primarily on static graphs, where all operations are executed on tensors with fixed shapes. For an LLM, all operations are executed with a fixed sequence length $L_{\text{max}}$ such that, given an input token sequence $x_{1},...,x_{L}$ with $L\leq L_{\text{max}}$, we operate on the padded sequence $x_{1},...,x_{L},x_{\text{pad}:L+1},...,x_{\text{pad}:L_{\text{max}}}$. Executing the prefill graph generates a single new token $x_{L+1}$ as well as key and value sequences $k^{l,h}_{1},...,k^{l,h}_{L},k^{l,h}_{\text{pad}:L+1},...,k^{l,h}_{\text{pad}:L_{\text{max}}}$ and $v^{l,h}_{1},...,v^{l,h}_{L},v^{l,h}_{\text{pad}:L+1},...,v^{l,h}_{\text{pad}:L_{\text{max}}}$ for the $l$th hidden layer and $h$th attention head.  [^1] For brevity, in future references to these variables we will omit the layer and head indices $l,h$. Similarly, we will only describe operations on the key values $k$, as our treatment of the value tensors $v$ is mostly identical. Executing the decode graph immediately after prefill generates new key and value tensors $k_{L+1}$ and $v_{L+1}$ that are used to update the KV cache in-place (i.e. the new output key sequence looks like $k_{1},...,k_{L},k_{L+1},k_{\text{pad}:L+2}...,k_{\text{pad}:L_{\text{max}}}$) and generates one more new token $x_{L+2}$. The new token and the updated KV cache are used as input for another call to the decode graph. This process is repeated until the EOS token or $x_{L_{\text{max}}}$ is generated. See Figure 3 (a) and (c) for a visualization of this process.
@@ 2.4 SnapKV
SnapKV is a KV cache compression method that mitigates the lossy compression seen in prior compression schemes by using a pooling-based clustering mechanism. The pooling layer helps to retain critical token clusters with higher attention weights, thus retaining the completeness of information. This clustering is particularly beneficial in retrieval tasks like needle-in-a-haystack, as we later show in the Retrieval subset of RULER in Table 1. Formally, given a prompt with length $L=L_{\text{prefix}}+L_{\text{obs}}$ divided into a prefix and an observation window, for the $l$th hidden layer we obtain the SnapKV index set $I_{l}$ by computing:
$C_{l}$ $=\sum_{i=0}^{L_{\text{obs}}}W_{l}[:,:,i,:]$ (1)
$I_{l}$ $=\text{Top}_{K}(\text{avgpool}(C_{l}))$ (2)
where $W_{l}\in\mathbb{R}^{B\times H\times L_{\text{obs}}\times L_{\text{prefix}}}$ is the softmax-normalized attention distribution of tokens in the observation window attending to prefix tokens for layer $l$ with batch size $B$ and number of attention heads HH. [^2] As with KV sequences, in future references we will omit the layer index $l$, as operations are identical across layers.
@@ 2.5 StreamingLLM
StreamingLLM xiao2024efficientstreaminglanguagemodels is a technique that enables LLMs to process sequences beyond their fixed context windows, effectively supporting unbounded streaming inputs and outputs. Conventional LLMs are limited by a fixed attention window, which constrains their ability to maintain coherence over long sequences. StreamingLLM addresses this by retaining a small set of attention sink tokens (the initial prompt tokens that consistently receive high attention across decoding steps) together with the most recent tokens. These tokens act as stable anchors that preserve long-term context, allowing the model to maintain continuity without storing the entire KV cache.
The core mechanism of StreamingLLM is a compact KV cache strategy that combines these sink and recent tokens to form a reduced attention context. During decoding, the model attends only to this subset, substantially lowering memory usage while preserving model accuracy.
@@ 3 Methodology
SnapStream is built on top of a static graph CB deployment by applying SnapKV to prefill KV caches and decoding with StreamingLLM. Following xiao2024efficientstreaminglanguagemodels, given an input sequence of length $L$, during prefill we extract the first $L_{\text{sink}}$ tokens and the last $L_{\text{recent}}$ tokens from the full KV cache. We apply SnapKV compression to the remaining $L-(L_{\text{sink}}+L_{\text{recent}})$ tokens, extracting the top $K$ tokens. We concatenate these three tensors along the sequence dimension to obtain a prefill KV cache of length $L_{\text{snapstream}}=L_{\text{sink}}+L_{\text{recent}}+K$. During decoding, we decode with a fixed KV cache size of $L_{\text{snapstream}}$, treating the middle $L_{\text{recent}}$ tokens as the rotating KV cache from StreamingLLM. See Figure 3(b) and (d) for a visualization of SnapStream prefill and decoding, respectively. In the following sections, we describe the layout of the SnapStream KV cache and the operation of SnapStream prefill and decoding in a production static-graph execution framework.
@@ 3.1.1 Case 1: $L<L_{sr}$
In this case, the key tensor cache is constructed during prefill as $k_{1},...,k_{L},k_{\text{pad}:L+1},..,k_{\text{pad}:L_{sr}},k_{\text{topk}:1},..,k_{\text{topk}:K}$. This structure is no different than that described in Section 2.3, as the Top-$K$ tokens $x_{\text{topk}:\cdot}$ are themselves padding tensors in this case. Accordingly, during decoding, the KV cache is updated in-place in the same manner as previously described in Section 2.3.
@@ 3.1.2 Case 2: $L\geq L_{sr}$
In this case, the key cache is constructed as $k_{1},...,k_{L_{\text{sink}}},k_{L-L_{\text{recent}}},...,k_{L},k_{\text{topk}:1},...,k_{\text{topk}:K}$. During prefill, we remove the middle $L-L_{sr}$ tokens from the KV cache, and extract the top $K$ key and value tensors from these tokens to popuate $k_{\text{topk}:1},...,k_{\text{topk}:K}$. During decoding, we maintain the same KV cache size by evicting the least recently used key tensor, $k_{L-L_{\text{recent}}}$, and replacing it with the new key tensor $k_{L+1}$. Note that in this case, the sink tensors and top-$K$ tensors are never modified during decoding.
@@ 3.2 Prefill
In order to reduce memory usage of the decoding stage, SnapStream applies SnapKV compression during prefill to compress an output KV cache along the sequence dimension. Adherent to the KV cache structure described in Section 3.1, we only apply SnapKV to compress the evicted tokens $k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}$. The sink token KV cache, recent token KV cache, and compressed SnapKV cache are then concatenated and transferred to the decoding node.
@@
  • We apply StreamingLLM after any positional encodings are applied to query, key, and value projections. This means that we do not apply positional encodings to cached KVs during decoding and use the original relative positional encodings. Although this may affect the ability of the model to generate beyond its trained attention window size as claimed in Section 3.2 of the StreamingLLM manuscript xiao2024efficientstreaminglanguagemodels, we believe that this change is well-motivated, as most modern open models are now trained to handle context lengths up to 128k grattafiori2024llama3herdmodels; deepseekai2025deepseekv3technicalreport; yang2025qwen3technicalreport, which is more than sufficient for most use cases.